OpenStudy (anonymous):

If the events A and B are independent, then P(A and B) = P(A)P(B). True or false. I say false..

OpenStudy (anonymous):

I believe, if there's some sort of cross dependence then it could be that P(A and B) is more likely than both happen to happen at the same time separately

OpenStudy (anonymous):

it is true.

OpenStudy (anonymous):

Damn I'm always failing this crap lol can you explain @surjithayer

OpenStudy (anonymous):

@surjithayer are the both terms the exact same thing ?

OpenStudy (anonymous):

if two events are independent then probability of one is independent of second and vice verse.

OpenStudy (anonymous):

so, because they're independent it doesn't matter if we ask P(A) and P(B) or P(A and B) because, P(A and B) is not more likely than P(A) and P(B)

OpenStudy (anonymous):

i dont understand independence :(

OpenStudy (anonymous):

I think understand it 0.o

OpenStudy (anonymous):

what is different, if there is dependence ? they're more likely to happen together right ?

OpenStudy (anonymous):

what confuses me is that P(A) and P(B) is also when they're together......

OpenStudy (anonymous):

Ohh yeah true... damn

OpenStudy (anonymous):

here I give one example. A die is thrown and 6 possible outcome are assumed to be equally likely. If E is the event :"the number appearing is a multiple of 3."and F The event:" the number appearing is even." Then show that E and F are independent. P(E)=P({3,6})=2/6=1/3 P(F)=P({2,4,6})=3/6=1/2 P(E and F)=P({6})=1/6 1/6=1/3 *1/2 P(E and F)=P(E) *P(F)